solve for (x,y): {(1/x + 1/y =6)/(1/x - 1/y= -4)
Answer by checkley77(12844)

(1/x + 1/y =6)
(1/x - 1/y = -4) subtract
---------------------------
2/y=10
10y=2
y=2/10
y=.2 ans.
(1/x + 1/y =6)
(1/x - 1/y = -4) add
------------------------
2/x=2
2x=2
x=1 ans.